use the given conditions to write an equation in the point-slope form and the slope-intercept form...thank you
slope=-2, passing through (0, -3)

Answer by venugopalramana(3286)

use the given conditions to write an equation in the point-slope form and the slope-intercept form...thank you
slope=-2 = M , passing through (0, -3) = X1,Y1
FORMULA IS
Y=M(X-X1)+Y1
Y=-2(X-0)+(-3)= -2X-3